# Thermodynamics of quantum processes: An operational framework for free energy and reversible athermality

### v1

@@ A.5 SDP for the max-free energy
A.6 Proof of Proposition 3
Proposition.
@@ A.7 Proof of Proposition 4
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{1-\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right).$ (156)
@@
$F_{\infty}^{\beta,\sqrt{\varepsilon}}[\mathcal{N}]$
$=\frac{1}{2\beta}D_{\infty}^{\sqrt{\varepsilon}}[\mathcal{N}\|\mathcal{T}^{\beta}]$ (160)
$\leq\frac{1}{2\beta}\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\infty}^{\sqrt{\varepsilon}}(\mathcal{N}(\psi)\|\mathcal{T}^{\beta}(\psi))$ (161)
$\leq\frac{1}{2\beta}\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\mathrm{H}}^{1-\varepsilon}(\mathcal{N}(\psi)\|\mathcal{T}^{\beta}(\psi))+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right)$ (162)
$=\frac{1}{2\beta}D_{\mathrm{H}}^{1-\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right)$ (163)
$=\frac{1}{2}F_{\mathrm{H}}^{\beta,1-\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right).$ (164)
@@ A.8 Free energy of a golden unit
From Proposition 3 we have
@@
To see that this value is the global minima for $F_{\infty}^{\beta}[\mathcal{U}]$ note that
$2^{\beta F_{\infty}^{\beta}[\mathcal{U}]}$ $=\operatorname{tr}{\operatorname{e}}^{\beta\widehat{H}_{A}}\operatorname{tr}^{-\beta\widehat{H}_{A}}$ (171)
@@
$=(\operatorname{tr}\mathbbm{1}_{A})^{2}$ (173)
$=|A|^{2}.$ (174)
Here we have used the Cauchy-Schwarz inequality for Hilbert-Schmidt inner product: $(\operatorname{tr}(A^{\dagger}B))^{2}\leq\operatorname{tr}(A^{\dagger}A)\operatorname{tr}(B^{\dagger}B)$. We conclude that $F_{\infty}^{\beta}[\mathcal{U}]\geq 2\beta^{-1}\ln|A|$ and the output Hamiltonian $\widehat{H}_{A}=c\mathbbm{1}_{A}$ gives the minimum free energy for unitary channels. ∎

### v2

@@ A.5 SDP for the max-free energy
A.6 Proof of Lemma 3
Lemma.
@@ A.7 Proof of Proposition 3
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{1-\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2}\ln\left(\frac{1}{1-\varepsilon}\right).$ (156)
@@
$D_{\infty}^{\sqrt{\varepsilon}}[\mathcal{N}\|\mathcal{T}^{\beta}]$ (160)
$\leq\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\infty}^{\sqrt{\varepsilon}}(\mathcal{N}(\psi)\|\mathcal{T}^{\beta}(\psi))$ (161)
$\leq\sup_{\psi\in\operatorname{St}(RA^{\prime})}D_{\mathrm{H}}^{1-\varepsilon}(\mathcal{N}(\psi)\|\mathcal{T}^{\beta}(\psi))+\ln\left(\frac{1}{1-\varepsilon}\right)$ (162)
$=D_{\mathrm{H}}^{1-\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]+\ln\left(\frac{1}{1-\varepsilon}\right).$ (163)
The proof follows since $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})=\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$ and $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})=\frac{1}{2}D_{\rm H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$. ∎
@@ A.8 Free energy of a golden unit
From Lemma 3 we have
@@
To see that this value is the global minimum for $F_{\infty}^{\beta}[\mathcal{U}]$, note that
${\rm e}^{\beta F_{\infty}^{\beta}[\mathcal{U}]}$ $=\operatorname{tr}{\operatorname{e}}^{\beta\widehat{H}_{A}}\operatorname{tr}\operatorname{e}^{-\beta\widehat{H}_{A}}$ (170)
@@
$=(\operatorname{tr}\mathbbm{1}_{A})^{2}=|A|^{2}.$ (172)
We have used the Cauchy-Schwarz inequality for Hilbert-Schmidt inner product: $(\operatorname{tr}(A^{\dagger}B))^{2}\leq\operatorname{tr}(A^{\dagger}A)\operatorname{tr}(B^{\dagger}B)$. We conclude that $F_{\infty}^{\beta}[\mathcal{U}]\geq 2\beta^{-1}\ln|A|$ and the output Hamiltonian $\widehat{H}_{A}=c\mathbbm{1}_{A}$ gives the minimum free energy for unitary channels. ∎